OpenStudy (salazarblack):

What is the last digit of number 7^2009

OpenStudy (salazarblack):

@dan815

OpenStudy (dan815):

There is a pattern for 7s

OpenStudy (dan815):

7, 49 , 243 , ... write a couple out see how it begins to repeat

OpenStudy (dan815):

There is more neater ways to do it, but you need to learn some modular arithmetic for it

OpenStudy (salazarblack):

isn't 7^343..

OpenStudy (salazarblack):

7^3 = 343*

OpenStudy (dan815):

okay xD

OpenStudy (dan815):

continue on with the pattern

OpenStudy (salazarblack):

till what point?

OpenStudy (dan815):

tll u see the last numbers repeating

OpenStudy (salazarblack):

okay did that.. xD

OpenStudy (dan815):

ok so tell me the pattern

OpenStudy (salazarblack):

7,49,343,2401,16807,117649,823543,5764801,40353607.... so that would be 7,9,3,1,7,9,3,1,7 and so on

OpenStudy (dan815):

very good now lets think, where would 2009 fall in this pattern

OpenStudy (dan815):

Do you have any ideas?

OpenStudy (dan815):

lets see the pattern of the last digits, **Only*** ~writing last digits~~~~ 7^1 = 7 7^2 = 9 7^3 = 3 7^4 = 1 7^5 = 7 7^6...

OpenStudy (dan815):

I ll let you think about it a little. Tell me what you come up with.

OpenStudy (dan815):

*Hint* every 4 exponents there is a repeating pattern*

OpenStudy (dan815):

think about that 4 and 2009, what can we do about this

OpenStudy (salazarblack):

i have an idea to sum all the last number until i get 2009 but that'll take some time xD

OpenStudy (salazarblack):

like 5*7^3+6*7^2+0*7^1 but this doesn't help much..

OpenStudy (dan815):

nope think simpler :)

OpenStudy (dan815):

for example ill give you a hint if i want to find like 2^8 lets say

OpenStudy (dan815):

i know every 4th exponent is 1 because its going 7,9,3,1, 7,9,3,1 7, 9, 3, 1

OpenStudy (dan815):

so like 7^8 = last number is 1 7^12 = last number is 1 again 7^16 = last number is 1

OpenStudy (dan815):

lets think about the closest multiple to 4 around 2009

OpenStudy (dan815):

so that we can work out the last numbers beside that number

OpenStudy (dan815):

for example since i know the pattern is 7 , 9 , 3, 1, 7,9,3,1 7,9,3,1 and if someone asked me what was 2^10 lets say I would find 2^8 which is 1, then i know 7 , 9 , 3 ,1 7, 9 ,3 ,1 ^--- I am there, so 2^10 i have to go up 2 more exponents and i get 9 for the last number.

OpenStudy (dan815):

So in same way for 2009 2008 is a multiple of 4 and 2^2008, last number is 1, so what would 2009 be?

OpenStudy (dan815):

oh my gosh ive been typing 2^2008 and 2^10.. excuse me for that lol it should be 7^2008 and 7^exponents..

OpenStudy (salazarblack):

it's okay xD

OpenStudy (salazarblack):

7

OpenStudy (dan815):

yep

OpenStudy (dan815):

Does it make sense?

OpenStudy (salazarblack):

it does but is there any other way to do this ...'cause this takes a lot of time and won't have much time on the test...

OpenStudy (salazarblack):

i won't*

OpenStudy (dan815):

hmm

OpenStudy (salazarblack):

+ i can't use calculator ...

OpenStudy (dan815):

oh really

OpenStudy (dan815):

Have you learnt mods?

OpenStudy (salazarblack):

i'm familiar with it but no..